# A Search for Classical Subsystems in Quantum Worlds

### v2

@@ VI Conclusions
In this work we have addressed the question: “Given a quantum world, determined by its energy spectrum, can we find factorizations of the global Hilbert space such that certain states of the system are robust to entanglement?”. The dynamical selection of such robust (pointer) states, known as einselection, has a key role in the emergence of classical from quantum allowing systems in pointer states to be tracked classically without consideration of various quantum processes. We have explored this question by constructing a cost function that is sensitive to the purity of the evolving subsystem quantum state. This gives us an operational tool to identify pointer states which we use to numerically explore systems comprised of several qubits. We support our numerical explorations with an analytic discussion of the various behaviors identified.
@@
The numerical results presented here are shown for a single qubit system and a two qubit environment ($n_{s}=1$, $n_{e}=2$). We explored a variety of sizes up to $n_{s}=2$ and $n_{e}=4$ and saw similar results. The analytical forms of the “destination” Hamiltonians that we derived in Section IV.4 exist for arbitrary $n_{s}$ and $n_{e}$ (with the exception of case (e) whose scaling behavior is unclear). For example, the block diagonal case is always available and would, in theory, give a vanishing cost function for any values of $n_{s}$ and $n_{e}$. Our algorithm was helpful in discovering what kinds of solutions can exist. The analytic forms were developed to understand the various solutions that the algorithm returned to us, and we view the analytic work as a key part of our results.
It would be interesting to extend this exploration to larger values of $n_{s}$ and $n_{e}$ and see what other kinds of solutions appear. Another important extension would involve expanding our treatment to numbers of subsystems larger than two. This is relevant to the Quantum Darwinism phenomena exhibited in our classical world [16] and also for constructing the sorts of probes we have mentioned in the context of applications to quantum engineering.
We have not attached a lot of conceptual importance to the fine details of our algorithm (despite the apparent need, in the case of the foundational questions for some sort of measure with which to compare the different realms). However, we are curious if our algorithm (or a related one) might be relevant for practical applications where probes of the qubit systems need to be suitably engineered. In principle the algorithm might reflect some aspects of the engineering process.
@@ VII Acknowledgements
Appendix A Block Diagonal Hamiltonians
@@
    $|0\rangle\rightarrow|00\rangle$ (25)
    $|1\rangle\rightarrow|01\rangle$
    $|2\rangle\rightarrow|10\rangle$
    $|3\rangle\rightarrow|11\rangle\,.$
@@
Appendix B Numerical Implementation
The optimization algorithm we employ is commonly used in the tensor network community. Favorable updates of the unitaries $A$ and $B$ with respect to the cost function in Eq. (9) are calculated using common density matrix renormalization group (DMRG) techniques [59] and the so-called Singular Value Decomposition (SVD) “trick”, which is related to finding the closest unitary matrix to a general matrix [60].
Figure 7: Violin plots capturing the statistics obtained by $100$ random initializations for each case. For each instance, the linear entropy of $\rho_{s}$ is evaluated for the set of training times (grey), as well as a set of times up to a maximum time (different shades of blue), and the highest entropy (i.e. worst-case) point from this set is collected. The various violins then depict the details of the statistics summarized in Fig. 2. Their interpretation is discussed extensively in Appendix C.
Figure 8: The figure depicts the linear entropy of the trained state (solid colors) for all of the classes of solutions discussed in Sec. IV.4. For comparison, we also show the evolution of the linear entropy of a random product state (dashed) in the fiducial factorization. All the solutions were obtained from the same fiducial Hamiltonian, the central spin (discussed in Sec. IV.2), with one central and three environment qubits. This figure describes the “quality” of each of the solutions (i.e. the depth of the minimum on which the algorithm converges) and should be compared with the last row of Fig. 7.
@@ Appendix B Numerical Implementation
To demonstrate the approach, assume a hypothetical cost function $\tilde{C}=1-\langle\psi_{1}|U_{1}U_{2}|\psi_{2}\rangle$ with fixed states $\psi_{1},\psi_{2}$ and trainable unitaries $U_{1},U_{2}$. It is known that the optimal (non-unitary) operator $O^{*}_{1}$ in the position of $U_{1}$ relative to all other components being fixed can be calculated via $O^{*}_{1}=|\psi_{1}\rangle\langle\psi_{2}|U^{\dagger}_{2}$. In other words, this is the operator that maximizes $\langle\psi_{1}|O_{1}U_{2}|\psi_{2}\rangle$ for any operator $O_{1}$. $O^{*}_{1}$ is commonly referred to as the environment tensor of $U_{1}$ and is calculated by multiplying (in tensor network language called “contracting”) all remaining terms after “removing” the term of interest (here $U_{1}$). The closest unitary $U^{*}_{1}$ to the non-unitary operator $O^{*}_{1}$ in terms of the Frobenius norm [61] can be calculated via the SVD, i.e., $O^{*}_{1}=\mathcal{U}\mathcal{S}\mathcal{V}^{\dagger}$ and $U^{*}_{1}=\mathcal{U}\mathcal{V}^{\dagger}$, where $\mathcal{U},\mathcal{V}^{\dagger}$ are orthogonal matrices and $\mathcal{S}$ a diagonal matrix carrying the singular values. We can now do the same optimization step for $U_{2}$ and continuously iterate over both unitaries to minimize the cost $\tilde{C}$. Just as with any optimizer in non-convex landscapes, there is no guarantee that different initializations of $U_{1}$ and $U_{2}$ all converge to a global optimum where $\tilde{C}=0$.
@@
Now consider the cost function in Eq. (10), which is not linear in the unitaries to be optimized, but is, in fact, quartic. In that case, the environment tensor that stems from removing one appearance of the unitaries denoted as $B$ is no longer the optimal operator in place of all appearances of $B$, but optimization can still be very successful in practice. The same approach can be taken for the unitary that prepares the initial input state (see Sec. IV.1). Optimization is usually improved by introducing a learning rate that interpolates between the former and proposed unitaries. We apply decaying learning rates for the unitaries, which add fractions of the current unitaries to the environment tensor before the SVD. This results in a smaller step in the space of unitaries into the direction of the approximately optimal unitary per iteration. We note that the particular optimization protocol may be fine-tuned to find global optima of the cost function more often, but this is not the focus of our work.
Appendix C Detailed Statistics
@@
The purpose of this numerical analysis is to inspect the distribution of solutions, i.e., how often we achieve solutions with certain properties with our particular optimization algorithm. We optimize 100 randomly initialized instances and record the maximum linear entropy of the reduced system density matrix until certain evolution times. We use two stopping criteria: either the instance achieves an “acceptance threshold” of $10^{-13}$ for the cost function Eq. (10), or runs for $10^{5}$ iterations, whichever comes first. This is why instances for which the former stopping criterion is invoked cluster around the $10^{-13}$ level (see the plots labeled ‘Training Times’ in e.g. the bottom row). These solutions also generalize well to late times. To see this, we show the maximum linear entropy until $t=10^{5.5}$. Note that we take $t=10^{5.5}$ as our “late time” limit, beyond which numerical instabilities can occur; if an instance performs well at this late time, we say that it “generalizes”. In Fig.8, we show that the decoupled, block diagonal, and furnace destination Hamiltonians discussed in Sec. IV.4 can be identified as these lowest entropy solutions.
@@
In the third row of Fig. 7, there is a class of solutions for which the maximum training time entropy forms a “blob” away from the acceptance threshold at the $\mathcal{O}(10^{-7})$ level. Interestingly, these solutions continue to generalize well in the late time limit with linear entropy much less than one, though much greater than $\mathcal{O}(10^{-12})$. As shown in in Fig. 8, the “Approximate Eigenstate” destination discussed in Sec. IV.4 can be identified with this class of solution which appears to form a local minimum in the cost landscape. For the random Hamiltonian, it is also clear from the widths of these “blobs”, relative to the width of the instances that cluster at the bottom, that these local minima solutions are just as numerous as those solutions which reach the acceptance threshold for the training times. However, in the weakly interacting regime of the central spin (interpolation points $0.99,0.999$), there are many more of these “blob” solutions with training cost $\approx\mathcal{O}(10^{-7})$. Given the description of the quality of each of the solutions in Fig. 8, one might be tempted to conclude that all these mid-grade solutions fall in the “Approximate Eigenstate” category. However, because this is a numerical exploration, it is possible to have solutions that are close to one of the other destination Hamiltonians since the algorithm may trigger the maximum iteration stopping criterion before it can converge to the global minimum. Thus, some of these mid-grade solutions may be approximately block diagonal or decoupled etc. We remind the reader again that this counting of the solutions is a reflection on the cost landscape and not (necessarily) a statement about a measure of physical significance (see discussion in Sec. V). Thus, we interpret that the over-abundance of the “blob” solutions in the weakly interacting limit of the central spin occurs due to the cost landscape getting overwhelmed by many local minima which prevent the algorithm from converging to one of the (possible many) global minima.
@@
As discussed in Sec. IV.4, there is also the “extended decoherence time” class of solutions for which the cost is much less than one for the training times but that they do not generalize well in the late time limit. These appear as the ‘neck” of the violins in the bottom row of Fig. 7 and their slow growth in entropy (relative to a random product state) is depicted in Fig. 8. Finally, a particularly intriguing case of the extended coherence time solutions can be seen in the middle row of Fig. 7. Here, the algorithm can achieve a low cost while optimizing the input state only on the fiducial system qubit. Such solutions are rare, and it is unclear whether this finding has a physical interpretation or whether identifying such behavior with a quasi-classical subsystem is like the case of “cloud watching” discussed at the end of Sec. IV.4.

### v3

@@ VI Conclusions
In this work we have addressed the question: Given a quantum world, determined by its energy spectrum, can we find factorizations of the global Hilbert space such that certain states of the system are robust to entanglement? The dynamical selection of such robust (pointer) states, known as einselection, has a key role in the emergence of classical from quantum allowing systems in pointer states to be tracked classically without consideration of various quantum processes. We have explored this question both numerically and analytically. On the numerical side, we introduced an algorithm to minimize a cost function that is sensitive to the purity of the evolving subsystem quantum state. This gives us an operational tool to identify pointer states which we use to numerically explore systems comprised of several qubits. Complementary to these numerical explorations, we have also found several analytic forms of Hamiltonian describing subsystems with quasi-classical properties. We find these analytic forms interesting in their own right, and also very helpful for interpreting the different solutions found by the algorithm.
@@
The numerical results presented here are shown for a single qubit system and a two qubit environment ($n_{s}=1$, $n_{e}=2$). We explored a variety of sizes up to $n_{s}=2$ and $n_{e}=4$ and saw similar results. The analytical forms of the “destination” Hamiltonians that we derived in Section IV.3 exist for arbitrary $n_{s}$ and $n_{e}$ (with the exception of case (e) whose scaling behavior is unclear). For example, the block diagonal case is always available and would, in theory, give a vanishing cost function for any values of $n_{s}$ and $n_{e}$. Our algorithm was helpful in discovering what kinds of solutions can exist. The analytic forms were developed to understand the various solutions that the algorithm returned to us, and we view the analytic work as a key part of our results.
It would be interesting to extend this exploration to larger values of $n_{s}$ and $n_{e}$ and see what other kinds of solutions appear. Another important extension would involve expanding our treatment to numbers of subsystems larger than two. This is relevant to the Quantum Darwinism phenomena exhibited in our classical world [zurek2009quantum-darwinism] and also for constructing the sorts of probes we have mentioned in the context of applications to quantum engineering. Furthermore, while we have demonstrated the existence of several “destinations” for some classes of Hamiltonians, an intriguing open question remains regarding the fundamental structures of the spectrum to which the realms owe their existence.
Finally, in this work we have focused on einselection as the tracer for the quantum to classical transition. In doing so, we have used predictability as the guiding principle for identifying quasi-classical subsystems. As briefly discussed in Sections I and III, einselection is but one aspect of classicality. We note that several authors (e.g. [Halliwell_1998_decoherent_hydrodynamics, Brun_Hartle_1999, Halliwell1999_decoherent_coarse_graining, Gell-Mann:2006ewf]) have considered the role of coarse graining, locally conserved quantities, and fast versus slow moving modes in determining which physical subsystems behave quasi-classically. We have not studied in detail how these broader ideas might relate to the “realms” we have identified. We regard these connections as an interesting topic for future work. We do note that behaviors we have identified based on the properties of the global energy eigenstates extend in straightforward ways to arbitrarily large systems and we expect that includes systems which admit these other analyses of classical phenomena (i.e. those in terms of coarse graining etc.).
We have not attached a lot of conceptual importance to the fine details of our algorithm. In Sec. V we considered foundational questions which, perhaps in a cosmological context, demand some sort of measure with which to give weight to the different coexisting realms. While our algorithm does operationally provide a measure, we have no reason to think it offers one that is relevant to such questions. On the other hand, we are curious if our algorithm (or a related one) might be relevant for practical applications where probes of the qubit systems need to be suitably engineered. In principle the algorithm might reflect some aspects of the engineering process.
@@ VII Acknowledgements
Appendix A Definitions
In this paper we use some words that are utilized differently by different sectors of our community. Here we clarify the usages we chose for this paper.
Subsystem
If a Hilbert space has a tensor product structure, e.g.
$\mathcal{H}=\mathcal{A}\otimes\mathcal{B}$ (25)
the factors $\mathcal{A}$ and $\mathcal{B}$ are candidates to be considered subsystems of the full quantum system. In practice it is helpful to consider additional properties (such as predictable behavior) that extend beyond this formal construction. That is the focus of this paper. Equivalent statements can be made for tensor product structures formed of any number of factors. We note that such constructions require that the dimension of $\mathcal{H}$ equals the product of the dimensions of its factors, which means not all Hilbert spaces can be factored in all possible ways.
Quantum coherence
Any system in a pure quantum state can be thought of as having quantum coherence. In practice quantum coherence gets more attention in some situations than others. For example a billiard ball in a pure wavepacket state (i.e. the Universe is in a product state of $|ball\rangle\otimes|rest\rangle$), is rarely explicitly regarded as quantum coherent, although if it is measured in a different basis quantum coherence would be manifest (i.e. through interference). The double slit diffraction pattern (which is typically measured in a completely different basis) is known for its quantum coherence. Further discussion along these lines can be found in [albrecht1994inflation-squeezing].
Decoherence
In the most general sense, decoherence is the loss of quantum coherence (as defined above) due to the entanglement with some other physical system. Many discussions focus on the case where the other system is a large “environment”, although that is not a requirement for a system to decohere. Some communities may use decoherence to refer to the onset of classicality or, more specifically, to refer to einselection (defined below) [zurek1993preferred-predictability-sieve]. In this paper, we use einselection to refer to a particular decohering phenomenon. There is also a conceptual and technical tool called “decohering histories”, which has only a loose connection to decoherence as defined here. We use the term “consistent histories” for that tool.
Einselection
or “environment induced superselection” refers to the case when interaction with the environment selects certain system states, called the “pointer states”, that remain stable (possibly only approximately) against entanglement [zurek1981pointer, zurek2003decoherence-and-the-quantum-origins, zurek2025decoherence].
To clarify the distinction we are making between general decoherence and decoherence that leads to einselection, we consider a quantum system consisting of two systems $A$ and $B$ which evolve according to a Hamiltonian $H$ which is a random and time independent Hermitian matrix in the full $A\otimes B$ space. For concreteness, let the factorization that partitions the two systems also be random. If the systems start in an initial product state then evolution according to the random $H$ will certainly entangle $A$ with $B$, i.e.
$e^{-iHt}|\psi\rangle_{A}|\phi\rangle_{B}=\sum_{i}\alpha_{i}(t)|a_{i}(t)\rangle|b_{i}(t)\rangle$ (26)
where the entangled state on the right is written in the Schmidt basis and there is more than one non-zero $\alpha_{i}(t)$. Consequently, the reduced system density matrix on system $A$ evolves as,
$\rho_{A}(t)=\sum_{k}\alpha_{k}^{2}(t)|a_{k}(t)\rangle\langle a_{k}(t)|\,$, (27)
where we use the fact that $\langle b_{i}(t)|b_{j}(t)\rangle=\delta_{ij}$ by the definition of the Schmidt basis. The two subsystems have decohered each other (Eq. (27) shows a mixed state). However, the random nature of $H$ means there will be no special system states $|\psi\rangle_{A}$ one could choose in Eq. (26) that would be robust to entanglement (as defined by Eq. (11)). Indeed, this claim is supported by our numerics in Fig. 1. Thus, despite there being decoherence, there are no pointer states, and so no einselection, for this example.
That this example is rather formal serves to highlight the fact that in the real world examples of decoherence and einselection are observed in parallel (hence the terms often being conflated). However, when conducting this sort of foundational study we have found the distinction between the two concepts to be helpful. And, indeed, it is einselection rather than decoherence that determines the emergence of subsystems.
Pointer States
States that are (possibly only approximately) robust to entanglement with the environment. Wavepackets of everyday classical objects have this property. This robustness also makes pointer states predictable since their evolution can be tracked without considering the effects of entanglement.
Predictability Sieve
An operational tool for determining the pointer states of a system by explicitly checking which states are the most robust to entanglement with the environment (and therefore behave predictably).
Branching
A term used to reference the different terms of an entangled wavefunction. In the relative state interpretation, einselection singles out the choice of basis with respect to which the different outcomes of each branch can be understood.
Realm
A factorization of the global Hilbert space into a system and an environment factor such that the system behaves quasi-classically. For this work, we have taken classicality to be associated with the existence of pointer states (see definition of Subsystem above). A given Hamiltonian can admit multiple realms, with no simple correspondences between the subsystems of the different realms.
Appendix B Block Diagonal Hamiltonians
@@
[…]
@@
Appendix C Numerical Implementation
C.1 Unitary optimization
The optimization algorithm we employ is commonly used in the tensor network community. It is based on constructing the cost function $\operatorname{Tr}[\rho^{2}]$ (see Eq. (12)) as one tensor network contraction, where the network here consists of several copies of the unitaries $A$ and $B$, $e^{-iHt}$ and $\psi\rangle_{w}$. By removing one of those tensors from the contraction, we compute its environment, whose transpose represents the ideal local update for the removed tensor. This approach is, for example, the preferred method for optimizing local tensors in the density matrix renormalization group (DMRG) algorithm [schollwock2005density]. We combine this with the so-called Singular Value Decomposition (SVD) “trick”, to find the closest unitary matrix to a general matrix [green1952orthogonal]. The implementation of our algorithm with efficient matrix operations using the JAX library can be found on Github: github.com/MSRudolph/subsystems.
@@ C.1 Unitary optimization
To demonstrate the approach, assume a hypothetical cost function $\tilde{C}=1-\langle\psi_{1}|U_{1}U_{2}|\psi_{2}\rangle$ with fixed states $\psi_{1},\psi_{2}$ and trainable unitaries $U_{1},U_{2}$. It is known that the optimal (non-unitary) operator $O^{*}_{1}$ in the position of $U_{1}$ relative to all other components being fixed can be calculated via $O^{*}_{1}=|\psi_{1}\rangle\langle\psi_{2}|U^{\dagger}_{2}$. In other words, this is the operator that maximizes $\langle\psi_{1}|O_{1}U_{2}|\psi_{2}\rangle$ for any operator $O_{1}$. $O^{*}_{1}$ is commonly referred to as the environment tensor of $U_{1}$ and is calculated by multiplying (in tensor network language called “contracting”) all remaining terms after “removing” the term of interest (here $U_{1}$). The closest unitary $U^{*}_{1}$ to the non-unitary operator $O^{*}_{1}$ in terms of the Frobenius norm [higham1989matrix] can be calculated via the SVD, i.e., $O^{*}_{1}=\mathcal{U}\mathcal{S}\mathcal{V}^{\dagger}$ and $U^{*}_{1}=\mathcal{U}\mathcal{V}^{\dagger}$, where $\mathcal{U},\mathcal{V}^{\dagger}$ are orthogonal matrices and $\mathcal{S}$ a diagonal matrix carrying the singular values. We can now do the same optimization step for $U_{2}$ and continuously iterate over both unitaries to minimize the cost $\tilde{C}$. Just as with any optimizer in non-convex landscapes, there is no guarantee that different initializations of $U_{1}$ and $U_{2}$ all converge to a global optimum where $\tilde{C}=0$.
@@
Now consider the cost function in Eq. (12), which is not linear in the unitaries to be optimized, but is, in fact, quartic. In that case, the environment tensor that stems from removing one appearance of the unitaries denoted as $B$ is no longer the optimal operator in place of all appearances of $B$, but optimization can still be very successful in practice. The same approach can be taken for the unitary that prepares the initial input state (see Sec. III.2). Optimization is usually improved by introducing a learning rate that interpolates between the former and proposed unitaries. We apply decaying learning rates for the unitaries, which add fractions of the current unitaries to the environment tensor before the SVD. This results in a smaller step in the space of unitaries into the direction of the approximately optimal unitary per iteration. We note that the particular optimization protocol may be fine-tuned to find global optima of the cost function more often, but this is not the focus of our work.
C.2 Training Times
In order for the cost function Eq. (12) to be faithful, the number of time-steps should be large (empirically, we find this to be $k\gtrsim 2d_{w}$ where $d_{w}\equiv\dim{(\mathcal{H}_{w}})$). This is because, for a small number of time steps, it is always possible to find specific states that minimize the cost at only the times on which the algorithm has been trained but that these states, unlike true pointer states, do not generalize to arbitrary times. This situation is explicitly depicted in the Fig. 6 (left panel). One can visualize this cost landscape as one that has many local minima for small values of $k$, but that many of these “evaporate” as $k$ increases, leaving behind solutions that generalize to times outside the training set.
Furthermore, for the faithfulness of the cost function, the set of training times needs to contain points sufficiently beyond the characteristic decoherence time of the Hamiltonian. This is important since the reduced density matrix of a generic initial product state will remain pure for $t\lesssim t_{\rm dec}$ where the decoherence time, $t_{\rm dec}$, is determined by $\lambda_{\rm max}$, the largest eigenvalue of the Hamiltonian, i.e. $t_{\rm dec}\equiv 1/\lambda_{\rm max}$. This effect of varying $T_{\rm train}$ on the generalizability of the solution can also be seen in Fig. 6 (right panel).
We note that it may be possible that if the ansatz for the unitary $B$ is chosen with care and in correspondence to the structure of the Hamiltonian, it is plausible that much fewer training times can be used. However, fully characterizing this algorithm is not the focus of this work.
Appendix D Detailed Statistics
Figure 7: Violin plots capturing the statistics obtained by $100$ random initializations for each case. For each instance, the linear entropy of $\rho_{s}$ is evaluated for the set of training times (grey), as well as a set of times up to a maximum time (different shades of blue), and the highest entropy (i.e. worst-case) point from this set is collected. The various violins then depict the details of the statistics summarized in Fig. 1. Their interpretation is discussed extensively in Appendix D.
Figure 8: The figure depicts the linear entropy of the trained state (solid colors) for all of the classes of solutions discussed in Sec. IV.3. For comparison, we also show the evolution of the linear entropy of a random product state (dashed) in the fiducial factorization. All the solutions were obtained from the same fiducial Hamiltonian, the central spin (discussed in Sec. IV.1), with one central and three environment qubits. This figure describes the “quality” of each of the solutions (i.e. the depth of the minimum on which the algorithm converges) and should be compared with the last row of Fig. 7.
@@ Appendix D Detailed Statistics
The purpose of this numerical analysis is to inspect the distribution of solutions, i.e., how often we achieve solutions with certain properties with our particular optimization algorithm. We optimize 100 randomly initialized instances and record the maximum linear entropy of the reduced system density matrix until certain evolution times. We use two stopping criteria: either the instance achieves an “acceptance threshold” of $10^{-13}$ for the cost function Eq. (12), or runs for $10^{5}$ iterations, whichever comes first. This is why instances for which the former stopping criterion is invoked cluster around the $10^{-13}$ level (see the plots labeled ‘Training Times’ in e.g. the bottom row). These solutions also generalize well to late times. To see this, we show the maximum linear entropy until $t=10^{5.5}$. Note that we take $t=10^{5.5}$ as our “late time” limit, beyond which numerical instabilities can occur; if an instance performs well at this late time, we say that it “generalizes”. In Fig.8, we show that the decoupled, block diagonal, and furnace destination Hamiltonians discussed in Sec. IV.3 can be identified as these lowest entropy solutions.
@@
In the third row of Fig. 7, there is a class of solutions for which the maximum training time entropy forms a “blob” away from the acceptance threshold at the $\mathcal{O}(10^{-7})$ level. Interestingly, these solutions continue to generalize well in the late time limit with linear entropy much less than one, though much greater than $\mathcal{O}(10^{-12})$. As shown in in Fig. 8, the “Approximate Eigenstate” destination discussed in Sec. IV.3 can be identified with this class of solution which appears to form a local minimum in the cost landscape. For the random Hamiltonian, it is also clear from the widths of these “blobs”, relative to the width of the instances that cluster at the bottom, that these local minima solutions are just as numerous as those solutions which reach the acceptance threshold for the training times. However, in the weakly interacting regime of the central spin (interpolation points $0.99,0.999$), there are many more of these “blob” solutions with training cost $\approx\mathcal{O}(10^{-7})$. Given the description of the quality of each of the solutions in Fig. 8, one might be tempted to conclude that all these mid-grade solutions fall in the “Approximate Eigenstate” category. However, because this is a numerical exploration, it is possible to have solutions that are close to one of the other destination Hamiltonians since the algorithm may trigger the maximum iteration stopping criterion before it can converge to the global minimum. Thus, some of these mid-grade solutions may be approximately block diagonal or decoupled etc. We remind the reader again that this counting of the solutions is a reflection on the cost landscape and not (necessarily) a statement about a measure of physical significance (see discussion in Sec. V). Thus, we interpret that the over-abundance of the “blob” solutions in the weakly interacting limit of the central spin occurs due to the cost landscape getting overwhelmed by many local minima which prevent the algorithm from converging to one of the (possible many) global minima.
@@
As discussed in Sec. IV.3, there is also the “extended decoherence time” class of solutions for which the cost is much less than one for the training times but that they do not generalize well in the late time limit. These appear as the ‘neck” of the violins in the bottom row of Fig. 7 and their slow growth in entropy (relative to a random product state) is depicted in Fig. 8. Finally, a particularly intriguing case of the extended coherence time solutions can be seen in the middle row of Fig. 7. Here, the algorithm can achieve a low cost while optimizing the input state only on the fiducial system qubit. Such solutions are rare, and it is unclear whether this finding has a physical interpretation or whether identifying such behavior with a quasi-classical subsystem perhaps akin to shapes found in the sky when cloud watching.